# Bound-electron self-energy calculations in Feynman and Coulomb gauges: detailed analysis

### v1

@@ II.2 Zero- and one-potential terms
where $\mathbf{\hat{p}}=\mathbf{p}/p$ and for a bound state $\widetilde{g}_{a}(p)=\widetilde{g}_{{n_{r}}_{a},\kappa_{a}}(p)$ and $\widetilde{f}_{a}(p)=\widetilde{f}_{{n_{r}}_{a},\kappa_{a}}(p)$. For the point-nucleus case, there are analytical expressions for $\widetilde{g}_{a}(p)$ and $\widetilde{f}_{a}(p)$, see, e.g., mohr_main. For an arbitrary potential, they can be obtained numerically according to:
@@
[…]
@@
where $z$ is the cosine of the angle between the vectors $\mathbf{p}$, $\mathbf{p}^{\prime}$, namely, $z=(\mathbf{\hat{p}}\cdot\mathbf{\hat{p}^{\prime}})$, $P_{l}$ is the Legendre polynomial. The final expression for the one-potential term reads as
@@
$\begin{gathered}X_{1}=A\widetilde{g}^{\prime}_{a}\widetilde{g}_{a}+\varepsilon_{a}(B_{1}+B_{2})K_{1}^{\prime}\widetilde{g}_{a}+\varepsilon_{a}(C_{1}+C_{2})\widetilde{g}^{\prime}_{a}K_{1}\\ +DK_{1}K_{1}^{\prime}+\varepsilon_{a}(H_{1}+H_{2})\widetilde{g}^{\prime}_{a}\widetilde{g}_{a}+G_{1}K_{1}^{\prime}\widetilde{g}_{a}+G_{2}\widetilde{g}^{\prime}_{a}K_{1},\\ X_{2}=A\widetilde{f}^{\prime}_{a}\widetilde{f}_{a}+\varepsilon_{a}(B_{1}+B_{2})K_{2}^{\prime}\widetilde{f}_{a}+\varepsilon_{a}(C_{1}+C_{2})\widetilde{f}^{\prime}_{a}K_{2}\\ +DK_{2}K_{2}^{\prime}-\varepsilon_{a}(H_{1}+H_{2})\widetilde{f}^{\prime}_{a}\widetilde{f}_{a}-G_{1}K_{2}^{\prime}\widetilde{f}_{a}-G_{2}\widetilde{f}^{\prime}_{a}K_{2},\\ K_{1}=\varepsilon_{a}\widetilde{g}_{a}+p\widetilde{f}_{a},\;K_{1}^{\prime}=\varepsilon_{a}\widetilde{g}^{\prime}_{a}+p^{\prime}\widetilde{f}^{\prime}_{a},\;K_{2}=\varepsilon_{a}\widetilde{f}_{a}+p\widetilde{g}_{a},\;K_{2}^{\prime}=\varepsilon_{a}\widetilde{f}^{\prime}_{a}+p^{\prime}\widetilde{g}^{\prime}_{a}.\end{gathered}$ (21)
For brevity, the dependence of wave functions on $p$ and $p^{\prime}$ is omitted. For the functions of $p^{\prime}$, an additional prime is added: $\widetilde{g}^{\prime}_{a}=\widetilde{g}_{a}(p^{\prime})$ and $\widetilde{f}^{\prime}_{a}=\widetilde{f}_{a}(p^{\prime})$.
@@ II.3 Many-potential term
As was said above, the many-potential contribution is calculated in coordinate space using the PW expansion. Therefore, let us start with the PW expansion of the bound-electron propagator (7), inspired by the explicit form of the Dirac-equation solution (2):
@@
where $\pi^{\pm\pm}(\mathbf{\hat{r}}_{1},\mathbf{\hat{r}}_{2})=\sum_{m_{n}}\chi_{\pm{\kappa_{n}},{m_{n}}}(\mathbf{\hat{r}}_{1})\chi_{\pm{\kappa_{n}},{m_{n}}}^{\dagger}(\mathbf{\hat{r}}_{2})$ and the components of the radial Green’s function $G_{\kappa_{n}}^{ij}$ are obtained in different ways depending on the method used. For fixed $\kappa_{n}$, the radial Dirac-equation spectrum is non-degenerate, so it is convenient to introduce the index $i_{n}$ that enumerates the solutions with the given angular symmetry. For example, for $G^{11}$ Eq. (7) results in
$G_{\kappa_{n}}^{11}(\omega,{r_{1},r_{2}})=\sum_{i_{n}}\frac{g_{{i_{n}},{\kappa_{n}}}(r_{1})g_{{i_{n}},{\kappa_{n}}}(r_{2})}{\omega-\varepsilon_{n}^{-}}.$ (23)
The index $i_{n}$ in Eq. (23) and similar expressions runs over both the positive- and negative-energy Dirac continua as well as all bound states. When using the DKB approach, the complete Dirac spectrum is replaced with the finite set of solutions of the equation (1). In this case, the index $i_{n}$ runs over this finite set. When using the GF approach, the radial Green’s function $G_{\kappa_{n}}^{ij}$ can be constructed from the homogeneous Dirac-equation solutions, bounded at zero, $\phi_{\kappa_{n}}^{0}=(g_{\kappa_{n}}^{0}\;\;f_{\kappa_{n}}^{0})^{\rm{T}}$, and at infinity, $\phi_{\kappa_{n}}^{\infty}=(g_{\kappa_{n}}^{\infty}\;\;f_{\kappa_{n}}^{\infty})^{\rm{T}}$, (see, e.g., Ref. sym12050800):
$G_{{\kappa_{n}}}(\omega,r_{1},r_{2})=\phi_{{\kappa_{n}}}^{\infty}(r_{1}){\phi_{{\kappa_{n}}}^{0}}^{\rm{T}}(r_{2})\theta(r_{1}-r_{2})+\phi_{{\kappa_{n}}}^{0}(r_{1}){\phi_{{\kappa_{n}}}^{\infty}}^{\rm{T}}(r_{2})\theta(r_{2}-r_{1}).$ (24)
Note that the solutions $\phi^{0}$ and $\phi^{\infty}$ are supposed to be normalized so that their Wronskian equals one. Expressions similar to Eqs. (22)-(24) can also be obtained for the free-electron Green’s function, which we will denote as $G^{(0)}$.
@@
In the GF case, because of the $\theta$-functions, radial integrations need to be handled carefully in Eq. (26), see, e.g., Ref. Art_trick. In the framework of the DKB approach, it is convenient to rewrite $G^{(2+)}$ in the form:
@@
where new states $|\check{n}\rangle$ are defined by:
$|\check{n}\rangle=\sum_{f}\frac{|f\rangle\langle f|V|n\rangle}{\omega-\varepsilon_{f}^{-}}\equiv{\begin{pmatrix}\check{g}_{n}(r)\chi_{\kappa_{n}m_{n}}(\mathbf{\hat{r}})\\ i\check{f}_{n}(r)\chi_{-\kappa_{n}m_{n}}(\mathbf{\hat{r}})\end{pmatrix}}$, (28)
and the multi-index $n$, as discussed above, corresponds to the complete set of quantum numbers: $n=(i_{n},\kappa_{n},m_{n})$. Then, using Eq. (28), the many-potential contribution is expressed as:
@@
The expression (29) can be readily adjusted to be used within the GF approach. For this aim, one need to reverse an analogue of Eq. (23) for $G^{(2+)}$:
$\sum_{i_{n}}\frac{\check{g}_{{i_{n}},{\kappa_{n}}}(r_{1})\check{g}_{{i_{n}},{\kappa_{n}}}(r_{2})}{\omega-\varepsilon_{n}^{-}}\rightarrow G^{(2+)11}_{\kappa_{n}}(\omega,r_{1},r_{2}).$ (30)
We use the following formula for the matrix element of the operator $I(\omega)$:
@@
where the reduced matrix elements on the right-hand side do not depend on the angular-momentum projections. Their explicit form in the Feynman and Coulomb gauges is given in Appendix C in Eqs. (71) and (72), respectively. The summation over the projections $m_{n}$ in Eq. (29) can be done analytically, and we obtain:
@@
II.4 Acceleration schemes
@@
The convergence of the PW-expansion series could be improved using different acceleration schemes. Their general idea is to subtract the term $\Delta\varepsilon_{a,x}^{\text{substr.}}$, which contains the slowly converging part of the PW expansion of the many-potential contribution. The index ‘‘$x$’’ indicates that the subtracted term should be calculated exactly in the same manner as Delta $\varepsilon_{a}^{\text{Mp}}$, that is in coordinate space using the PW expansion. The same term, evaluated separately but not necessarily within the same approach, so we denote it as $\Delta\varepsilon_{a}^{\text{substr.}}$ without the additional index ‘‘$x$’’, should then be added back. This idea can be illustrated by the formula:
$\Delta\varepsilon_{a}^{\text{Mp}}\rightarrow(\Delta\varepsilon_{a}^{\text{Mp}}-\Delta\varepsilon_{a,x}^{\text{substr.}})+\Delta\varepsilon_{a}^{\text{substr.}}.$ (33)
@@
It is implied that the PW-expansion series for the two contributions in the brackets in Eq. (33) are subtracted term by term and only then extrapolated. The resulting difference is expected to demonstrate a better convergence than the initial series for the many-potential contribution. Therefore, If one can calculate $\Delta\varepsilon_{a}^{\text{substr.}}$ with high accuracy, it will determine the overall improvement in the calculations of the SE correction.
@@
Within the two-potential acceleration scheme (see the Introduction section), the subtracted term is the two-potential contribution, $\Delta\varepsilon_{a,x}^{\text{substr.}}=\Delta\varepsilon_{a,x}^{2\text{p}}$. When adding this term, one must evaluate it with high precision in order to benefit from applying the scheme. A natural idea is to perform such calculations in momentum space without using the PW expansions. However, this problem is complicated and leads to complex multidimensional integrals, see, e.g., Ref. Yer2010, where a similar approach was employed in the point-nucleus case for the SE correction to the hyperfine splitting. For this reason, an alternative was proposed in Refs. PhysRevLett.98.173004; Art_trick, where the difference between the many- and two-potential terms was addressed using the FBS method, while the two-potential term itself was also considered in coordinate space but employing the Green’s functions method and truncating the corresponding PW summations at higher values of angular momenta. We note that in this case, it is not necessary to solve the system of differential equations numerically for an arbitrary potential, since the PW expansion of the free-electron propagator can be expressed in terms of the spherical Bessel functions and the desired two-potential term can be obtained by combining three such propagators. In the present work, we study this two-potential scheme but for simplicity the difference between the many- and two-potential terms is also calculated based on the Green’s function approach. We refer to this term as the three-plus-potential contribution.
The Sapirstein-Cheng (SC) scheme sap_orig is based on the idea that the dominant contribution from the free-electron propagators comes from the region where both spatial arguments are close to each other. Therefore, one can obtain an approximation to the two-potential term by ‘‘transposing’’ the potentials from the inner electron line to the vertices, where the photon propagator is attached, see Fig. 3. We refer to this approximation as the quasi-two-potential term and denote it as $\Delta\varepsilon_{a,x}^{\text{substr.}}=\Delta\tilde{\varepsilon}_{a,x}^{2\text{p}}$.
@@
A solid advantage of this scheme is the fact that it is fairly easy to calculate the quasi-two-potential contribution with high accuracy in momentum space. We denote the result as $\Delta\tilde{\varepsilon}_{a,p}^{2\text{p}}$ and use it as $\Delta\varepsilon_{a}^{\text{substr.}}$. Let us discuss the SC scheme in more detail.
In the coordinate-space part of the calculations, we should construct the PW expansion
$\widetilde{G}^{(3+)}=G^{(2+)}-\sum_{f}\frac{|Vf\rangle\langle fV|}{(\varepsilon_{a}-\omega-\varepsilon_{f}^{-})^{3}},$ (34)
which provides an approximation for the contribution involving three or more interactions with the binding potential. Here $|Vf\rangle=V|f\rangle$. Having this expansion, one can directly evaluate the expression $\Delta\tilde{\varepsilon}_{a,x}^{3+}=\Delta\varepsilon_{a}^{\text{Mp}}-\Delta\tilde{\varepsilon}_{a,x}^{2\text{p}}$, which we will refer as the quasi-three-plus-potential contribution. Then, it remains to calculate the $\Delta\tilde{\varepsilon}_{a,p}^{2\text{p}}$ in momentum space. The potentials that have been transposed into vertices can be formally taken into account by multiplying the wave function of the external state $|a\rangle$ by them. An issue arises of calculating the Fourier transform of such products. Since $V$ is assumed to be spherically symmetric, this can be done by a straightforward generalization of Eqs. (13)-(15). We represent the result in the following form:
@@
which readily follows from the spectral representation for $G^{(0)}$ and the orthogonality of the Dirac basis, one obtains
@@
where $\mathrm{p}=(p_{0},\mathbf{p})$. The explicit form of the derivatives of the free-electron self-energy operator $\Sigma_{R}^{(0)}$ in both gauges is given in Appendix D. After performing the angular integration, which is carried out similarly to that in the zero-potential contribution, one can obtain:

### v2

@@ II.2 Zero- and one-potential terms
where $\mathbf{\hat{p}}=\mathbf{p}/p$, $l_{a}=|\kappa_{a}+1/2|-1/2$, and for the bound state $\widetilde{g}_{a}(p)=\widetilde{g}_{{n_{r}}_{a}\kappa_{a}}(p)$ and $\widetilde{f}_{a}(p)=\widetilde{f}_{{n_{r}}_{a}\kappa_{a}}(p)$. In the point-nucleus case, there are analytical expressions for $\widetilde{g}_{a}(p)$ and $\widetilde{f}_{a}(p)$, which can be found, e.g., Ref. Mohr et al. (1998). For an arbitrary potential, they can be obtained numerically according to:
@@
where $j_{l}$ is the spherical Bessel function of the first kind.
@@
where $z$ is the cosine of the angle between the vectors $\mathbf{p}$ and $\mathbf{p}^{\prime}$, namely, $z=(\mathbf{\hat{p}}\cdot\mathbf{\hat{p}^{\prime}})$, and $P_{l}$ is the Legendre polynomial. The final expression for the one-potential term reads as
@@
$X_{1}$ $=A\widetilde{g}^{\prime}_{a}\widetilde{g}_{a}+\varepsilon_{a}(B_{1}+B_{2})K_{1}^{\prime}\widetilde{g}_{a}+\varepsilon_{a}(C_{1}+C_{2})\widetilde{g}^{\prime}_{a}K_{1}$
$+DK_{1}K_{1}^{\prime}+\varepsilon_{a}(H_{1}+H_{2})\widetilde{g}^{\prime}_{a}\widetilde{g}_{a}+G_{1}K_{1}^{\prime}\widetilde{g}_{a}+G_{2}\widetilde{g}^{\prime}_{a}K_{1},$ (21)
$X_{2}$ $=A\widetilde{f}^{\prime}_{a}\widetilde{f}_{a}+\varepsilon_{a}(B_{1}+B_{2})K_{2}^{\prime}\widetilde{f}_{a}+\varepsilon_{a}(C_{1}+C_{2})\widetilde{f}^{\prime}_{a}K_{2}$
$+DK_{2}K_{2}^{\prime}-\varepsilon_{a}(H_{1}+H_{2})\widetilde{f}^{\prime}_{a}\widetilde{f}_{a}-G_{1}K_{2}^{\prime}\widetilde{f}_{a}-G_{2}\widetilde{f}^{\prime}_{a}K_{2},$ (22)
where
$K_{1}=\varepsilon_{a}\widetilde{g}_{a}+p\widetilde{f}_{a},\;K_{1}^{\prime}=\varepsilon_{a}\widetilde{g}^{\prime}_{a}+p^{\prime}\widetilde{f}^{\prime}_{a},\;K_{2}=\varepsilon_{a}\widetilde{f}_{a}+p\widetilde{g}_{a},\;K_{2}^{\prime}=\varepsilon_{a}\widetilde{f}^{\prime}_{a}+p^{\prime}\widetilde{g}^{\prime}_{a}.$ (23)
For brevity, the dependence of wave functions on $p$ and $p^{\prime}$ is omitted. For the functions of $p^{\prime}$, an additional prime is added. Thus, the shorthand notations are $\widetilde{g}_{a}=\widetilde{g}_{a}(p)$, $\widetilde{f}_{a}=\widetilde{f}_{a}(p)$, $\widetilde{g}^{\prime}_{a}=\widetilde{g}_{a}(p^{\prime})$, and $\widetilde{f}^{\prime}_{a}=\widetilde{f}_{a}(p^{\prime})$.
@@ II.3 Many-potential term
As noted above, the remaining many-potential contribution is calculated in coordinate space using the PW expansion. Therefore, let us start with the PW expansion of the bound-electron propagator (7), inspired by the explicit form of the Dirac-equation solution (2):
@@
where $\pi^{\pm\pm}(\mathbf{\hat{r}}_{1},\mathbf{\hat{r}}_{2})=\sum_{m_{n}}\chi_{\pm{\kappa_{n}}{m_{n}}}(\mathbf{\hat{r}}_{1})\chi_{\pm{\kappa_{n}}{m_{n}}}^{\dagger}(\mathbf{\hat{r}}_{2})$ and the components of the radial Green’s function $G_{\kappa_{n}}^{ij}$ are obtained in different ways depending on the method used: FBS or GF.
For a fixed $\kappa_{n}$, the spectrum of the radial Dirac equation is non-degenerate, so it is convenient to introduce the index $i_{n}$ that enumerates solutions for the given angular symmetry:
$\begin{cases}\left(-\dfrac{d}{dr}+\dfrac{\kappa_{n}-1}{r}\right)f_{{i_{n}}{\kappa_{n}}}+\left(V+m_{e}\right)g_{{i_{n}}{\kappa_{n}}}=\varepsilon_{{i_{n}}{\kappa_{n}}}g_{{i_{n}}{\kappa_{n}}}\,,\\ \left(\dfrac{d}{dr}+\dfrac{\kappa_{n}+1}{r}\right)g_{{i_{n}}{\kappa_{n}}}+\left(V-m_{e}\right)f_{{i_{n}},{\kappa_{n}}}=\varepsilon_{{i_{n}}{\kappa_{n}}}f_{{i_{n}}{\kappa_{n}}}\,.\end{cases}$ (25)
Then, Eq. (7) results for, e.g., $G^{11}$ in
$G_{\kappa_{n}}^{11}(\omega,{r_{1},r_{2}})=\sum_{i_{n}}\frac{g_{{i_{n}}{\kappa_{n}}}(r_{1})g_{{i_{n}}{\kappa_{n}}}(r_{2})}{\omega-\varepsilon_{{i_{n}}{\kappa_{n}}}^{-}}.$ (26)
We stress that the index $i_{n}$ in Eq. (26) and similar expressions runs over both the positive- and negative-energy Dirac continua as well as all bound states for the chosen $\kappa_{n}$. When using the DKB approach, for each $\kappa_{n}$ the spectrum of the radial Dirac equation (25) is replaced with the finite set of solutions. In this case, the index $i_{n}$ runs over these finite sets. When using the GF approach, the radial Green’s function $G_{\kappa_{n}}^{ij}$ can be constructed from the solutions of the homogeneous Dirac equation, which are bounded at zero, $\phi_{\kappa_{n}}^{0}=(g_{\kappa_{n}}^{0}\;\;f_{\kappa_{n}}^{0})^{\text{{T}}}$, and at infinity, $\phi_{\kappa_{n}}^{\infty}=(g_{\kappa_{n}}^{\infty}\;\;f_{\kappa_{n}}^{\infty})^{\text{{T}}}$, (see, e.g., Ref. Yerokhin and Maiorova (2020)):
$G_{{\kappa_{n}}}(\omega,r_{1},r_{2})=\phi_{{\kappa_{n}}}^{\infty}(r_{1}){\phi_{{\kappa_{n}}}^{0}}^{\text{{T}}}(r_{2})\theta(r_{1}-r_{2})+\phi_{{\kappa_{n}}}^{0}(r_{1}){\phi_{{\kappa_{n}}}^{\infty}}^{\text{{T}}}(r_{2})\theta(r_{2}-r_{1}).$ (27)
Note that the solutions $\phi^{0}$ and $\phi^{\infty}$ are supposed here to be normalized so that their Wronskian equals one. Expressions similar to Eqs. (24)-(27) can also be obtained for the free-electron Green’s function, which we will denote as $G^{(0)}$.
@@
In the GF case, because of the $\theta$-functions, radial integrations need to be handled carefully in Eq. (29), see, e.g., Ref. Artemyev et al. (2013). In the framework of the DKB approach, it is convenient to rewrite $G^{(2+)}$ in the form:
@@
where the new states $|\check{n}\rangle$ are defined by:
$|\check{n}\rangle=\sum_{f}\frac{|f\rangle\langle f|V|n\rangle}{\omega-\varepsilon_{f}^{-}}\equiv{\begin{pmatrix}\check{g}_{i_{n}\kappa_{n}}(r)\chi_{\kappa_{n}m_{n}}(\mathbf{\hat{r}})\\ i\check{f}_{i_{n}\kappa_{n}}(r)\chi_{-\kappa_{n}m_{n}}(\mathbf{\hat{r}})\end{pmatrix}}$. (31)
Then, using Eq. (31), the many-potential contribution is expressed as:
@@
The expression (32) can be readily adjusted to be used within the GF approach. For this aim, one needs to reverse an expression similar to Eq. (26):
$\sum_{i_{n}}\frac{\check{g}_{{i_{n}}{\kappa_{n}}}(r_{1})\check{g}_{{i_{n}}{\kappa_{n}}}(r_{2})}{\omega-\varepsilon_{n}^{-}}\rightarrow G^{(2+)11}_{\kappa_{n}}(\omega,r_{1},r_{2}).$ (33)
The matrix element of the operator $I(\omega)$ can be expressed as follows:
@@
[…]
@@
II.4 Partial-wave expansion acceleration schemes
@@
The convergence of the PW-expansion series could be improved using different acceleration schemes. Their general idea is to separate out and subtract the term $\Delta\varepsilon_{a,x}^{\text{subtr}}$, which contains the slowly converging part of the PW expansion of the many-potential contribution. The index ‘‘$x$’’ in $\Delta\varepsilon_{a,x}^{\text{subtr}}$ indicates that the subtracted term should be calculated exactly in the same manner as $\Delta\varepsilon_{a}^{\text{Mp}}$, that is in coordinate space using the PW expansion. Naturally that the same term, evaluated separately, but not necessarily within the same approach, so we denote it as $\Delta\varepsilon_{a}^{\text{subtr}}$ without the additional index ‘‘$x$’’, should then be added back. The discussed idea can be illustrated by the formula:
$\Delta\varepsilon_{a}^{\text{Mp}}\rightarrow(\Delta\varepsilon_{a}^{\text{Mp}}-\Delta\varepsilon_{a,x}^{\text{subtr}})+\Delta\varepsilon_{a}^{\text{subtr}}.$ (36)
@@
It is implied that the PW-expansion series for the two contributions in the brackets in Eq. (36) are subtracted term by term and only then extrapolated in $k$. The resulting difference is expected to demonstrate a better convergence (at least, in terms of absolute values) than the initial series for the many-potential contribution. Therefore, if one can calculate $\Delta\varepsilon_{a}^{\text{subtr}}$ with high accuracy, it will determine the overall improvement in the calculations of the many-potential contribution to the SE correction.
@@
Within the two-potential acceleration scheme (see the Introduction section), the subtracted term is the two-potential contribution, $\Delta\varepsilon_{a,x}^{\text{subtr}}=\Delta\varepsilon_{a,x}^{2\text{p}}$. When adding this term back, one must evaluate it with high precision in order to benefit from applying the scheme. A natural idea is to perform the corresponding calculations in momentum space without using the PW expansions. However, this problem is complicated and leads to complex multidimensional integrals, see, e.g., Ref. Yerokhin and Jentschura (2010b), where a similar approach was employed in the point-nucleus case for the SE correction to the hyperfine splitting. For this reason, an alternative was proposed in Refs. Artemyev et al. (2007, 2013), where the difference between the many- and two-potential terms was addressed using the FBS method, while the two-potential term itself was also considered in coordinate space but employing the GF method and truncating the corresponding PW summations at higher values of angular momenta. We note that in this case, it is not necessary to solve the system of differential equations numerically for an arbitrary potential, since the PW expansion of the free-electron propagator can be expressed in terms of the spherical Bessel functions Mohr et al. (1998) and the desired two-potential term can be obtained by combining three such propagators. In the present work, we study this two-potential scheme but for simplicity the difference between the many- and two-potential terms is also treated based on the GF approach. The PW expansion for this difference arises from the Green’s function
$G^{(3+)}=G^{(2+)}-\sum_{f_{1},f_{2},f_{3}}\frac{|f_{1}\rangle\langle f_{1}|V|f_{2}\rangle\langle f_{2}|V|f_{3}\rangle\langle f_{3}|}{(\omega-\varepsilon_{f_{1}}^{-})(\omega-\varepsilon_{f_{2}}^{-})(\omega-\varepsilon_{f_{3}}^{-})}\,.$ (37)
We refer to the term corresponding to $G^{(3+)}$ as the three-plus-potential contribution, $\Delta\varepsilon_{a}^{\text{(3+)p}}\equiv(\Delta\varepsilon_{a}^{\text{Mp}}-\Delta\varepsilon_{a,x}^{2\text{p}})$.
@@
The Sapirstein-Cheng (SC) scheme Sapirstein and Cheng (2023) is based on the idea that the dominant contribution from the free-electron propagators $G^{(0)}(\omega,\mathbf{r}_{1},\mathbf{r}_{2})$ comes from the region where both their spatial arguments are close to each other Mohr (1974b), $\mathbf{r}_{1}\approx\mathbf{r}_{2}$. Therefore, one can obtain an approximation to the two-potential term by ‘‘transposing’’ the potentials from the inner electron line to the vertices, where the photon propagator is attached, see Fig. 3. We refer to this approximation as the quasi-two-potential term and denote it as $\Delta\varepsilon_{a,x}^{\text{subtr}}=\Delta\tilde{\varepsilon}_{a,x}^{2\text{p}}$. The difference between the many- and quasi-two-potential terms, $\Delta\tilde{\varepsilon}_{a,x}^{(3+)\text{p}}\equiv\Delta\varepsilon_{a}^{\text{Mp}}-\Delta\tilde{\varepsilon}_{a,x}^{2\text{p}}$, referred to as the quasi-three-plus-potential contribution, can be studied in the coordinate space using the PW expansion for the Green’s function
$\widetilde{G}^{(3+)}=G^{(2+)}-\sum_{f}\frac{|Vf\rangle\langle fV|}{(\varepsilon_{a}-\omega-\varepsilon_{f}^{-})^{3}}$ (38)
Here, $|Vf\rangle=V|f\rangle$. A solid advantage of the SC scheme is the fact that it is fairly easy to calculate the quasi-two-potential contribution with high accuracy in momentum space. We denote the result as $\Delta\tilde{\varepsilon}_{a,p}^{2\text{p}}$ and use it as $\Delta\varepsilon_{a}^{\text{subtr}}$. The potentials that have been transposed into vertices can be formally taken into account by multiplying the wave function of the external state $|a\rangle$ by them. An issue arises of calculating the Fourier transform of such products. Since $V$ is assumed to be spherically symmetric, this can be done by a straightforward generalization of Eqs. (13)-(15). We represent the result in the following form:
@@
one obtains
@@
where $\mathrm{p}=(p_{0},\mathbf{p})$. The explicit form of the derivatives of the free-electron self-energy operator $\Sigma_{R}^{(0)}$ with respect to $p_{0}$ in both gauges is given in Appendix D. After performing the angular integration, which is carried out similarly to that in the zero-potential contribution, one can obtain: